The answer is not 889,000. Please correctly solve for taxable income and tax liability. Thank you! I will leave a thumbs up, if correct. Asset Building Sparkle Corporation stock Sales Price Cost $670,000 $653,000 229,000 281,000 Accumulated Depreciation $31,000
Answer #1

In case of sale of Real property, Corporate is subject to Section 291 capture. According to Section 291, corporate taxpayers to recapture 20% of the lessor of the gain realized or accumulated depreciation taken.
The recaptured gain is taxed as ordinary gain.
The remaining gain is taxed as Section 1231 gain.

1)a Realized gain on Building = $670,000 - ($653,000 - $31,000) = $48,000

Section 291 recapture - 20% of lessor or realized gain or Acc. dep. ($31,000 * 20%) = $6,200 ordinary gain

Section 1231 gain = $48,000 - $6,200 = $41,800 (1231 gain)

1)b Realized loss on sale of Corp stock = $229,000 - $281,000 = $52,000 LTCL

Section 1231 gain will be offset against the loss on sale of stock and the rest loss can be carry back up to 3 years or carry forward up to 5 years.

1) Taxable income after transactions = Taxable income before transactions + Ordinary Income = $893,000 + $6,200 = $899,200

2) Tax Liability = $899,200 * 21% = $188,832

Please up vote. Thanks!